OpenStudy (anonymous):

lim x->0 cos(theta)tan(theta)/theta

OpenStudy (anonymous):

\[\Large \lim_{x \to 0} \frac{\cos (x) \tan (x)}{x}? \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ \Large \lim_{x \to 0} \cos (x) \cdot \underbrace{\frac{\tan x}{x}}_{ \to 1}\]

OpenStudy (anonymous):

but the x is theta

OpenStudy (anonymous):

doesn't matter for this kind of problem.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just notice the identity \[\Large \lim_{x \to 0} \frac{\cos x \tan x}{x}= \lim_{x \to 0} \cos x \lim_{x \to 0} \frac{\tan x}{x} \] and know that \[\lim_{x \to 0 } \frac{\tan x}{x}=1 \]

OpenStudy (anonymous):

so it would be